OpenStudy (anonymous):

Check my work please? 1. (2x+5) (3x+7) I got : 35x^3+35 2. (-8a^3 b^4)^2 I got: 64a^6 b^8

OpenStudy (johnweldon1993):

not quite....try it again....remember 2x * 3x = ? 2x * 7 = ? 5 * 3x = ? 5 * 7 = ? same thing for the second one

OpenStudy (whpalmer4):

First is incorrect, second is correct.

OpenStudy (whpalmer4):

For the first one, remember that you have to multiply each component of the second term by each component of the first term. \[(a+b)(c+d) = a*c + a*d + b*c + b*d\]

OpenStudy (anonymous):

wouldnt I just combine like terms?

OpenStudy (anonymous):

I got 6x+14x+15x+35

OpenStudy (whpalmer4):

\[2x*3x \ne 6x\]

OpenStudy (anonymous):

still a bit confused

OpenStudy (anonymous):

6x^2 ?

OpenStudy (anonymous):

6x^2+39x +35 is that right?

OpenStudy (whpalmer4):

Getting closer... :-)

OpenStudy (johnweldon1993):

VERY close...1 more time

OpenStudy (whpalmer4):

\[(2x+5)(3x+7) = 2x*3x + 2x*7 + 5*3x + 5*7 = \]

OpenStudy (anonymous):

39x^2

OpenStudy (whpalmer4):

\[2x*3x = \]\[2x*7=\]\[5*3x=\]\[5*7=35\] (I did the last one for you :-))

OpenStudy (johnweldon1993):

lets do it the step by step 2x * 3x is? 2x * 7 is? 5 * 3x is? 5 * 7 is?

OpenStudy (anonymous):

6x^2......14x.....15x.....35

OpenStudy (johnweldon1993):

RIGHT so 6x² + 14x + 15x + 35 what is your final answer?

OpenStudy (anonymous):

6x^2 +29x+35

OpenStudy (johnweldon1993):

There you go! :)

OpenStudy (anonymous):

i feel so supid when it comes to math lol. thanks both of you

OpenStudy (johnweldon1993):

lol, just try to break questions down into steps...makes it a lot easier :) and no problem!

OpenStudy (whpalmer4):

Yeah, simple steps, done methodically, attention to detail...